# Online Convex Optimization with Memory and Limited Predictions

### v1

@@ Abstract
We study the problem of online convex optimization with memory and predictions over a horizon $T$. At each time step, a decision maker is given some limited predictions of the cost functions from a finite window of future time steps, i.e., values of the cost function at certain decision points in the future. The decision maker then chooses an action and incurs a cost given by a convex function that depends on the actions chosen in the past. We propose an algorithm to solve this problem and show that the dynamic regret of the algorithm decays exponentially with the prediction window length. Our algorithm contains two general subroutines that work for wider classes of problems. The first subroutine can solve general online convex optimization with memory and bandit feedback with $\sqrt{T}$-dynamic regret with respect to $T$. The second subroutine is a zeroth-order method that can be used to solve general convex optimization problems with a linear convergence rate that matches the best achievable rate of first-order methods for convex optimization. The key to our algorithm design and analysis is the use of truncated Gaussian smoothing when querying the decision points for obtaining the predictions. We complement our theoretical results using numerical experiments.
@@ 1 Introduction
Online decision making is a ubiquitous framework that captures a wide range of real-world applications, such as machine learning, control systems and finance [1]. In this framework, a decision maker interacts with an unknown environment by sequentially choosing a decision point $x_{t}$ at each time step and incurs a cost $f_{t}(x_{t})$ given by some underlying cost function $f_{t}(\cdot)$, and the goal is to minimize the accumulative cost over a horizon $t=1,\dots,T$. For convex cost functions $f_{t}(\cdot)$, the problem is known as Online Convex Optimization (OCO) which has been widely studied in the past decades (see, e.g., [2, 3] for comprehensive survey). The OCO framework may be further split into two settings. In the full information setting, the decision maker receives complete knowledge of the cost function $f_{t}(\cdot)$ after choosing $x_{t}$ for time step $t$ [4]; in the bandit information setting, the decision maker only receives $f_{t}(x_{t})$ (and potentially $f_{t}(\bar{x}_{t})$ at some perturbed points $\bar{x}_{t}$ from $x_{t}$) after choosing $x_{t}$ for time step $t$ [5, 6, 7]. Since the decision maker needs to choose the decision points in an online manner (i.e., before knowing all the cost functions), regret is a typical performance metric used for algorithms proposed to solve OCO, which compares the accumulative cost incurred by the algorithm to the minimum achievable cost in hindsight [4]. For non-stationary environment [8], the minimum achievable cost typically requires time-varying decisions over the horizon $T$ and the corresponding regret is referred to as dynamic regret [9, 10, 11].
@@
In many online decision making problems, the cost incurred at the current time step depends not only on the decision made for the current time step but also on the decisions made in the past. For example, when the underlying environment is characterized by a dynamical system whose state evolves according to the chosen actions, the cost incurred in the current step is governed by the current action and system state which in turn depends on past actions and system states. Such a framework of OCO (also known as OCO with memory) has been studied in [12, 13, 14] and specialized to the online control problem of linear dynamical system [15, 16, 17, 18, 19]. Due to the long-term dependency of the cost on the decision history, i.e., the current decision also influences cost at future steps, there is a line of research that also considers predictions of the cost in a window of future time steps [20, 21, 22, 23, 24]. We refer to this framework as OCO with memory and predictions. However, the works mentioned above have certain limitations. In [21, 22], the cost is assumed to be only a function of decisions in the current step and one step in the past. The classical control algorithm, model predictive control (MPC) [25, 26, 27], considers the receding horizon optimization for the next $W$ stages. However, similar to [24, 23], the predictions of the full information of the cost functions are assumed to be available in the MPC algorithms.
Motivated by these limitations in the existing works, we study the problem of OCO with memory and limited (i.e., partial) predictions. Specifically, we consider the scenario that at each time step, the decision maker is allowed to only query a limited number of points and obtain potentially noisy predictions of the cost functions from a finite-length prediction window at these points, and the incurred cost has a long-term dependency on the past decisions. Studying the partial prediction setting is natural since it indeed corresponds to the partial (i.e., bandit) information setting that has been widely studied in the classic OCO framework (without memory or prediction) as we described above. Our approaches and contributions are summarized as follows.
@@
$\bullet$ Algorithm Design and Regret Analysis. In Section 3, we propose an algorithm (Algorithm 1) to solve the problem of OCO with memory and limited predictions. Our algorithm extends that in [23] for full information setting to the partial information setting when receiving the predictions. We prove in Section 4 that the algorithm enjoys a dynamic regret that decays exponentially with the length of the prediction window and matches the dynamic regret of the algorithm in [23]. Our algorithm contains two general subroutines that can be applied to solve wider classes of problems as we describe next.
@@
$\bullet$ OCO with Memory and Bandit feedback. The first subroutine of Algorithm 1 can be applied to the general OCO problem with memory and bandit feedback. We show in Theorem 1 our algorithm achieves a dynamic regret of $\tilde{\mathcal{O}}(\sqrt{T})$, where $\mathcal{O}(\cdot)$ hides logarithmic factors in $T$. Our dynamic regret matches that of the algorithm proposed in [8] for OCO with memory and full information of cost functions (up to logarithmic factors in $T$).
@@
$\bullet$ Faster Convergence Rate of Zeroth-Order Method. The second subroutine of Algorithm 1 can be viewed as a zeroth-order (i.e., derivative-free) method for solving general (offline) convex minimization problems. Zeroth-order methods have typically been used to solve convex minimization problems when one only has access to a function value oracle while computing the derivative of the function is not allowed [28, 29]. For smooth and strongly convex cost functions, our algorithm achieves the same linear convergence rate as the first-order method using gradient descent [30, 31], which improves the convergence rate of the zeroth-order method proposed in [28]. We achieve this speedup by introducing a novel truncated Gaussian smoothing (formally introduced in Definition 1) when choosing the query points for the function value oracle, i.e., we perturb the decision point $x_{t}$ at time step $t$ by truncated Gaussian random vectors and query the perturbed points to obtain predictions of the function values.
$\bullet$ Noisy Function Value Oracle. Algorithm 1 and results in Theorems 1-2 hold for the setting when there is noise in the function value oracle used in Algorithm 1 to obtain the predictions. Such a noisy model subsumes the exact function value oracle as a special case.
@@
More Related Work. First, Bandit OCO and zeroth-order methods for convex minimization are naturally related to each other since in both scenarios, a general recursive way to update the action has the form: $x_{t+1}=x_{t}-\alpha_{t}g_{t}$, where $g_{t}$ is obtained from the function value oracle and can be viewed as an estimate of the true gradient at $x_{t}$. A major difference is that the performance of algorithms for bandit OCO is characterized by regret while the performance of algorithms for zeroth-order methods is characterized by convergence rate. However, it has been shown that an algorithm for bandit OCO with regret guarantee can be converted to an algorithm for zeroth-order methods with convergence guarantee [32, 29]. In this work, we propose an algorithm that combines a bandit OCO algorithm and a zeroth-order algorithm to solve the OCO problem with memory and limited predictions. Second, our algorithm design and analysis rely on the use of truncated Gaussian smoothing. The methods of Gaussian smoothing or smoothing over a sphere have been used in bandit OCO or zeroth-order methods [5, 30, 33, 34]. We show that the truncated Gaussian smoothing inherits the desired properties of Gaussian smoothing while also allowing us to prove a faster convergence rate of the zeroth-order method for smooth and strongly convex cost functions. Finally, several extensions of the framework in this paper can be considered, including extending to OCO with unbounded memory (i.e., the current cost depends on all the past decisions) [35] and nonsmooth or nonconvex cost functions [36].
@@
Notation and terminology. The sets of integers and real numbers are denoted as $\mathbb{Z}$ and $\mathbb{R}$, respectively. For a set $\mathcal{X}$, $\mathcal{X}^{n}$ denotes the corresponding Cartesian product $\mathcal{X}\times\cdots\times\mathcal{X}$. For a real number $a$, let $\lfloor a\rfloor$ be the largest integer that is smaller than or equal to $a$. For a vector $x\in\mathbb{R}^{n}$, let $\left\|x\right\|=\sqrt{x^{\top}x}$ be its euclidean norm. Let $\left\|P\right\|_{F}=\sqrt{\operatorname{\mathrm{tr}}(PP^{\top})}$ be the Frobenius norm of $P\in\mathbb{R}^{n\times m}$. Let $I_{n}$ be an $n$-dimensional identity matrix. Given any integer $n\geq 1$, $[n]\triangleq\{1,\dots,n\}$.
@@ 2 Problem Formulation
We consider an Online Convex Optimization (OCO) framework over $T\in\mathbb{Z}_{\geq 1}$ time steps. At each $t\in[T]$, a decision maker (i.e., agent) chooses an action $x_{t}\in\mathcal{X}\subseteq\mathbb{R}^{d}$ and incurs a cost $f_{t}(x_{t-h+1},\dots,x_{t})$, where $f_{t}:\mathbb{R}^{dh}\to\mathbb{R}$ is convex and $h\in\mathbb{Z}_{\geq 1}$. In other words, the cost incurred at any time step $t\in\mathbb{Z}_{\geq 1}$ has memory that depends on the current action $x_{t}$ and the past actions $x_{t-1},x_{t-2},\dots,x_{t-h+1}$. Denote
@@
Even though solving (3) requires complete knowledge of $f_{1}(\cdot),\dots,f_{T}(\cdot)$ a priori, in practice one may have access to predictions of $f_{t}(\cdot),\cdots,f_{t+W-1}(\cdot)$ for some $W\in\mathbb{Z}_{\geq 1}$ when choosing $x_{t}\in\mathcal{X}$ at time step $t\in\mathbb{Z}_{\geq 1}$ [21, 22, 26, 27]. However, the available predictions can be inaccurate and scarce (i.e., limited). Unlike existing works as discussed in the introduction, in this paper, we focus on a general framework to capture the inaccuracy and scarcity of the predictions. Specifically, at each time step $t\in\mathbb{Z}_{\geq 1}$, the predictions can be obtained by first running sequences of actions $\{(x_{t}^{j},\dots,x_{t+W-1}^{j})\}_{j=1}^{p}$ on a simulator and then receiving the corresponding costs $\{l_{t}(x_{t-h+1:t}^{j}),\dots,l_{t+W-1}(x_{t+W-h:t+W-1}^{j})\}_{j=1}^{p}$, where $l_{k}:\mathcal{X}^{h}\to\mathbb{R}$ for all $k\in\{t,\dots,t+W-1\}$. In other words, the simulator is associated with a model $l_{t}(\cdot)$ that describes the true cost $f_{t}(\cdot)$ with some model misspecification and can be viewed as a noisy oracle of the function values of $f_{t}(\cdot)$. Based on the available information, the agent then chooses an action $x_{t}\in\mathcal{X}$. Note that for each time step $t\in\mathbb{Z}_{\geq 1}$, one may obtain multiple cost prediction trajectories from the simulator by running multiple sequences of actions on the simulator in parallel (or running a single trajectory on multiple simulators). To summarize, the agent does the following:
  • At the beginning of time step $t\in\mathbb{Z}_{\geq 1}$, run sequences of actions $\{(x_{t}^{j},\dots,x^{j}_{t+W-1})\}_{j=1}^{p}$ on simulator $l_{t}(\cdot)$ and receive the predictions $\{(l_{t}(x^{j}_{t-h+1:t}),\dots,l_{t+W-1}(x^{j}_{t+W-h:t+W-1})\}_{j=1}^{p}$.
@@
[…]
Remark 1.
Based on the above arguments, we see that the information setting considered in our framework corresponds to the bandit information in the classic OCO framework (without memory or predictions) [5, 7, 37, 38, 11], since only values of the cost function are available when the agent makes decisions.
@@ 3 Algorithm Design
To tackle the limited predictions (i.e., bandit information) in our problem setting, we adopt the zeroth-order (i.e., derivative-free) method that has been used for both bandit OCO and (offline) convex optimization problems (e.g., [5, 28]). To this end, we extend the algorithm design in [24, 20] for full information setting to bandit information setting. The detailed steps are summarized in Algorithm 1. Using similar arguments to those in [23], one can argue that the overall algorithm can be viewed as using a zeroth-order method for solving (2), which consists of two phases: initialization and zeroth-order method update. Specifically, denoting $\bm{x}^{j}\triangleq(x_{1}^{j},\dots,x_{T}^{j})$ for any $j\in\{0,\dots,K-1\}$, the algorithm first initializes $\bm{x}^{0}$ using an OCO scheme, and then iteratively computes the update $\bm{x}^{j+1}=\Pi_{\mathcal{X}}(\bm{x}^{j}-\alpha\bm{g}^{j})$ for $j=0,\dots,K-1$, where $\Pi_{\mathcal{X}}(\cdot)$ denotes the projection onto the convex set $\mathcal{X}^{T}$ and $\bm{g}^{j}=(\bm{g}^{j}_{1},\dots,\bm{g}^{j}_{T})\in\mathbb{R}^{dT}$ with $\bm{g}^{j}_{s}=\sum_{k=s}^{s+h-1}g_{k}^{j}$ for all $s\in[T]$. In particular, the term $\bm{g}^{j}$ is obtained from $l_{1}(\cdot),\dots,l_{T}(\cdot)$ queried at certain points (see lines 6 and 16 of Algorithm 1) which can be viewed as an estimate of the true gradient $\nabla C_{T}(\bm{x}^{j})$. Note that we simply set $f_{t}(\bm{x})=l_{t}(\bm{x})=0$ for all $t\leq 0$, all $t>T$ and all $\bm{x}\in\mathcal{X}^{h}$ in Algorithm 1, and recall that we have set $x_{t}^{0}=\bar{x}_{0}\in\mathcal{X}$ for all $t\leq 1$. We summarize several key features of Algorithm 1.

### v3

@@ Abstract
This paper addresses an online convex optimization problem where the cost function at each step depends on a history of past decisions (i.e., memory), and the decision maker has access to limited predictions of future cost values within a finite window. The goal is to design an algorithm that minimizes the dynamic regret against the optimal sequence of decisions in hindsight. To this end, we propose a novel predictive algorithm and establish strong theoretical guarantees for its performance. We show that the algorithm’s dynamic regret decays exponentially with the length of the prediction window. Our algorithm comprises two general subroutines of independent interest. The first subroutine solves online convex optimization with memory and bandit feedback, achieving a $\sqrt{TV_{T}}$-dynamic regret, where $V_{T}$ measures the variation of the optimal decision sequence. The second is a zeroth-order method that attains a linear convergence rate for general convex optimization, matching the best achievable rate of first-order methods. The key to our algorithm is a novel truncated Gaussian smoothing technique when querying the decision points to obtain the predictions. We validate our theoretical results with numerical experiments.
@@ 1 Introduction
Online decision making is a ubiquitous framework that captures a wide range of real-world applications, such as machine learning, control systems, and finance cesa2006prediction. In this framework, a decision maker interacts with an unknown environment by sequentially choosing a decision point $x_{t}$ at each time step and incurs a cost $f_{t}(x_{t})$ given by some underlying cost function $f_{t}(\cdot)$, and the goal is to minimize the accumulative cost over a horizon $t=1,\dots,T$. For convex cost functions $f_{t}(\cdot)$, the problem is known as Online Convex Optimization (OCO), which has been widely studied in the past decades (see, e.g., hazan2016introduction; shalev2012online for a comprehensive survey). The OCO framework may be further split into two settings. In the full information setting, the decision maker receives complete knowledge of the cost function $f_{t}(\cdot)$ after choosing $x_{t}$ for time step $t$ zinkevich2003online; in the bandit information setting, the decision maker only receives $f_{t}(x_{t})$ (and potentially $f_{t}(\bar{x}_{t})$ at some perturbed points $\bar{x}_{t}$ from $x_{t}$) after choosing $x_{t}$ for time step $t$ flaxman2005online; agarwal2010optimal; saha2011improved. Since the decision maker needs to choose the decision points in an online manner (i.e., before knowing all the cost functions), regret is a typical performance metric used for algorithms proposed to solve OCO, which compares the accumulative cost incurred by the algorithm to the minimum achievable cost in hindsight zinkevich2003online. For the non-stationary environment zhao2022non, the minimum achievable cost typically requires time-varying decisions over the horizon $T$, leading to dynamic regret yang2016tracking; jadbabaie2015online; zhao2021bandit.
@@
In many online decision-making problems, the cost incurred at the current time step depends not only on the decision made for the current time step but also on the decisions made in the past. For example, when the underlying environment is characterized by a dynamical system whose state evolves according to the chosen actions, the cost incurred at the current step is governed by the current action and system state, which in turn depends on past actions and system states. Such a framework of OCO (i.e., OCO with memory) has been studied in arora2012online; anava2015online; zhou2023efficient and specialized to the online control problem of linear dynamical systems agarwal2019online; simchowitz2020improper; yan2023online; ye2024online; ye2024learning. Due to the long-term dependency of the cost on the decision history, i.e., the current decision also influences cost at future steps, there is a line of research that also considers predictions of the cost in a window of future time steps lin2012online; chen2015online; chen2016using; li2020online; li2019control. We refer to this framework as OCO with memory and predictions. However, the works mentioned above have certain limitations. In chen2015online; chen2016using, the cost is assumed to be only a function of decisions in the current step and one step in the past. The classical control algorithm, model predictive control (MPC) rawlings2012postface; zeilinger2011real; angeli2016theoretical, considers the receding horizon optimization for the next $W$ stages. However, similar to li2019control; li2020online, the predictions of the full information of the cost functions are assumed to be available in the MPC algorithms.
Motivated by these limitations in the existing works, we study the problem of OCO with memory and limited (i.e., partial) predictions. At each time step, the decision maker is allowed to only query a limited number of points and obtain potentially noisy predictions of the cost functions from a finite-length prediction window at these points, and the incurred cost has a long-term dependency on the past decisions. Focusing on this partial (or bandit) prediction setting is crucial as it reflects many practical applications where the full cost function or its gradient is unavailable. For instance, when interacting with a complex physical system, a proprietary software, or a black-box simulator, it is often impossible or computationally prohibitive to determine the underlying mathematical model nesterov2017random; balasubramanian2022zeroth. Instead, the decision-maker can only perform queries by inputting a set or sequence of actions and observing the resulting cost. Such “query-and-observe” interaction precisely corresponds to the bandit feedback model studied in classic OCO problems flaxman2005online; hazan2014bandit. We summarize our approaches and contributions.
@@
$\bullet$ Algorithm Design and Regret Analysis. In Section 3, we give an algorithm (Algorithm 1) to solve the problem of OCO with memory and limited predictions. Our algorithm extends that in li2020online for the full information setting to the partial information setting when making the predictions. We prove in Section 4 that the algorithm enjoys a dynamic regret that decays exponentially with the length of the prediction window and matches the dynamic regret of the algorithm in li2020online. Our algorithm contains two general subroutines that can be applied to solve wider classes of problems as we describe next.
@@
$\bullet$ OCO with Memory and Bandit feedback. The first subroutine of Algorithm 1 can be applied to general OCO with memory and bandit feedback. We show in Theorem 1 our algorithm achieves a $\sqrt{TV_{T}}$-dynamic regret, where $V_{T}$ is the variation length of the optimal solution. Our dynamic regret matches that of the algorithm in zhao2022non for OCO with memory and full information of cost functions (up to logarithmic factors in $T$).
@@
$\bullet$ Faster Convergence Rate of Zeroth-Order Method. Our core contribution lies in the second subroutine of Algorithm 1, which can be viewed as a zeroth-order (i.e., derivative-free) method for solving general (offline) convex minimization problems. For smooth and strongly convex cost functions, our algorithm achieves the same linear convergence rate as the first-order method using gradient descent nesterov2013introductory; bubeck2015convex, which improves the convergence rate of the zeroth-order method proposed in nesterov2017random. In particular, the convergence rate of our method is independent of the problem horizon length $T$, while applying the method in nesterov2017random yields a convergence rate that scales as $\mathcal{O}(1/T)$. We achieve this speedup by introducing a novel truncated Gaussian smoothing (formally introduced in Definition 1) when choosing the query points for the function value oracle, i.e., we perturb the decision point $x_{t}$ at time step $t$ by truncated Gaussian random vectors and query the perturbed points to obtain predictions of the function values.
@@
More Related Work. First, Bandit OCO and zeroth-order methods for convex minimization are naturally related to each other since in both scenarios, a general recursive way to update the action has the form: $x_{t+1}=x_{t}-\alpha_{t}g_{t}$, where $g_{t}$ is obtained from the function value oracle and can be viewed as an estimate of the true gradient at $x_{t}$. A major difference is that the performance of algorithms for bandit OCO is characterized by regret while the performance of algorithms for zeroth-order methods is characterized by convergence rate. However, it has been shown that an algorithm for bandit OCO with regret guarantee can be converted to an algorithm for zeroth-order methods with convergence guarantee shamir2013complexity; shamir2017optimal. In this work, we propose an algorithm that combines a bandit OCO algorithm and a zeroth-order algorithm to solve the OCO problem with memory and limited predictions. Second, our algorithm design and analysis rely on the use of truncated Gaussian smoothing. The methods of Gaussian smoothing or smoothing over a sphere have been used in bandit OCO or zeroth-order methods flaxman2005online; nesterov2013introductory; berahas2022theoretical; balasubramanian2022zeroth. We show that the truncated Gaussian smoothing inherits the desired properties of Gaussian smoothing while also allowing us to prove a faster convergence rate of the zeroth-order method for smooth and strongly convex cost functions. Finally, several extensions of the framework in this paper can be considered, including extending to OCO with unbounded memory (i.e., the current cost depends on all the past decisions) kumar2024online and nonsmooth or nonconvex cost functions lin2022gradient.
@@
Notation and terminology. The sets of integers and real numbers are denoted as $\mathbb{Z}$ and $\mathbb{R}$, respectively. For a set $\mathcal{X}$, $\mathcal{X}^{n}$ denotes the corresponding Cartesian product $\mathcal{X}\times\cdots\times\mathcal{X}$. For a real number $a$, let $\lfloor a\rfloor$ be the largest integer that is smaller than or equal to $a$. For a vector $x\in\mathbb{R}^{n}$, let $\left\|x\right\|=\sqrt{x^{\top}x}$ be its euclidean norm. Let $\left\|P\right\|_{F}=\sqrt{\operatorname{\mathrm{tr}}(PP^{\top})}$ be the Frobenius norm of $P\in\mathbb{R}^{n\times m}$. Let $I_{n}$ be an $n$-dimensional identity matrix. Given any integer $n\geq 1$, $n]\triangleq\{1,\dots,n\}$.
@@ 2 Problem Formulation
We consider an Online Convex Optimization (OCO) framework over $T\in\mathbb{Z}_{\geq 1}$ time steps. At each $t\in[T]\triangleq\{1,\dots,T\}$, a decision maker (i.e., agent) chooses an action $x_{t}\in\mathcal{X}\subseteq\mathbb{R}^{d}$ and incurs a cost $f_{t}(x_{t-h+1},\dots,x_{t})$, where $f_{t}:\mathbb{R}^{dh}\to\mathbb{R}$ is convex and $h\in\mathbb{Z}_{\geq 1}$. In other words, the cost incurred at any time step $t\in\mathbb{Z}_{\geq 1}$ has memory that depends on the current action $x_{t}$ and the past actions $x_{t-1},x_{t-2},\dots,x_{t-h+1}$. This “cost with memory” structure models a broad class of problems where decisions have long-term effects. Common application scenarios include control systems, where a sequence of actions affects future system states due to the dynamics of the system li2019control. The memory parameter $h$ is an intrinsic characteristic of the problem environment, representing the duration for which past actions remain influential. In practice, its value can be determined by domain knowledge of the specific system or estimated from historical data zhao2022non; zhou2023efficient. To proceed, denote
@@
Even though solving (3) requires complete knowledge of $f_{1}(\cdot),\dots,f_{T}(\cdot)$ a priori, in practice one may have access to predictions of $f_{t}(\cdot),\cdots,f_{t+W-1}(\cdot)$ for some $W\in\mathbb{Z}_{\geq 1}$ when choosing $x_{t}\in\mathcal{X}$ at time step $t\in\mathbb{Z}_{\geq 1}$ chen2015online; chen2016using; zeilinger2011real; angeli2016theoretical. However, the available predictions can be inaccurate and scarce (i.e., limited). Unlike existing works as discussed in the introduction, in this paper, we focus on a general framework to capture the inaccuracy and scarcity of the predictions. Specifically, at each time step $t\in\mathbb{Z}_{\geq 1}$, the predictions can be obtained by first running sequences of actions $\{(x_{t}^{j},\dots,x_{t+W-1}^{j})\}_{j=1}^{p}$ on a simulator and then receiving the corresponding costs $\{l_{t}(x_{t-h+1:t}^{j}),\dots,l_{t+W-1}(x_{t+W-h:t+W-1}^{j})\}_{j=1}^{p}$, where $l_{k}:\mathcal{X}^{h}\to\mathbb{R}$ for all $k\in\{t,\dots,t+W-1\}$. In other words, the simulator is associated with a model $l_{t}(\cdot)$ that describes the true cost $f_{t}(\cdot)$ with some model misspecification and can be viewed as a noisy oracle of the function values of $f_{t}(\cdot)$. Based on the available information, the agent then chooses an action $x_{t}\in\mathcal{X}$. Note that for each time step $t\in\mathbb{Z}_{\geq 1}$, one may obtain multiple cost prediction trajectories from the simulator by running multiple sequences of actions on the simulator in parallel (or running a single trajectory on multiple simulators). This prediction model, which allows querying future costs for chosen action sequences, is designed to represent practical scenarios where a high-fidelity simulator or a parallel testing environment is available. For example, in robotics, a planner can query a physics simulator to evaluate the cost of different trajectories, or in energy systems, an operator can simulate various generation schedules to predict their impact before commitment. This type of “query access” to a value-based oracle corresponds to the widely-studied bandit feedback setting in the classic OCO framework (without memory or predictions) flaxman2005online; saha2011improved; hazan2014bandit; hazan2016optimal; zhao2021bandit, since only values of the cost function are available when the agent makes decisions. Our work applies this well-established and practical model to the specific context of OCO with memory and predictions. To summarize, the agent does the following in each time step $t\geq 1$:
  • At the beginning of time step $t\geq 1$, run sequences of actions $\{(x_{t}^{j},\dots,x^{j}_{t+W-1})\}_{j=1}^{p}$ on simulator $l_{t}(\cdot)$ and receive the predictions $\{(l_{t}(x^{j}_{t-h+1:t}),\dots,$ $l_{t+W-1}(x^{j}_{t+W-h:t+W-1})\}_{j=1}^{p}$.
@@
  • Incur the cost $f_{t}(x_{t-h+1:t})$ and proceed to the next time step t+1t+1.111We assume that the predictions at time step $t$ can be obtained before the algorithm enters the next time step.
@@ 3 Algorithm Design
To tackle the limited predictions (i.e., bandit information) in our problem setting, we adopt the zeroth-order (i.e., derivative-free) method that has been used for both bandit OCO and (offline) convex optimization problems (e.g., flaxman2005online; nesterov2017random). To this end, we extend the algorithm design in li2019control; lin2012online for the full information setting to the bandit information setting. The detailed steps are summarized in Algorithm 1. Using similar arguments to those in li2020online, one can argue that the overall algorithm can be viewed as using a zeroth-order method for solving (2), which consists of two phases: Initialization (lines 3-7) and zeroth-order method update (lines 8-20). Specifically, let us denote $\bm{x}^{j}\triangleq(x_{1}^{j},\dots,x_{T}^{j})$ for any $j\in[0,\dots,K-1]$. In the first phase, Algorithm 1 uses an OCO subroutine to generate an initial solution sequence $\bm{x}^{0}=(x_{1}^{0},\dots,x_{T}^{0})$ that is competitive to the optimal solution $\bm{x}^{\star}$ given by (2), i.e., this sequence serves as a high-quality “warm-start” for the subsequent steps of Algorithm 1. In the second phase, Algorithm 1 employs a zeroth-order method to iteratively compute the update $\bm{x}^{j+1}=\Pi_{(1-\xi^{\prime})\mathcal{X}}(\bm{x}^{j}-\alpha\bm{g}^{j})$ for $j=0,\dots,K-1$, where $\Pi_{(1-\xi^{\prime})\mathcal{X}}(\cdot)$ denotes the projection onto set $(1-\xi^{\prime})\mathcal{X}^{T}$ and $\bm{g}^{j}=(\bm{g}^{j}_{1},\dots,\bm{g}^{j}_{T})\in\mathbb{R}^{dT}$ with $\bm{g}^{j}_{s}=\sum_{k=s}^{s+h-1}g_{k}^{j}$ for all $s\in[T]$. In particular, the term $\bm{g}^{j}$ is obtained from $l_{1}(\cdot),\dots,l_{T}(\cdot)$ queried at certain points (see lines 6 and 16 of Algorithm 1) which can be viewed as an estimate of the true gradient $\nabla C_{T}(\bm{x}^{j})$. Note that we simply set $f_{t}(\bm{x})=l_{t}(\bm{x})=0$ for all $t\leq 0$, all $t>T$ and all $\bm{x}\in\mathcal{X}^{h}$ in Algorithm 1, and recall that we have set $x_{t}^{0}=\bar{x}_{0}\in\mathcal{X}$ for all $t\leq 1$. We summarize several key features of Algorithm 1.
Feasibility of Queried and Action Points. In lines 7 and 17 of Algorithm 1, we project onto smaller sets $(1-\xi)\mathcal{X}$ and $(1-\xi^{\prime})\mathcal{X}$ compared to the original feasible set $\mathcal{X}$, where $\xi,\xi^{\prime}$ are input parameters to the algorithm. Later in Section 4, we will show that via proper choices of input parameters $\delta,\delta^{\prime},\xi,\xi^{\prime}$ to Algorithm 1, projecting onto these smaller sets ensures that both the action points $x_{t}^{K}$ (line 21) and the perturbed points $\bar{x}_{r}^{0},\tilde{x}_{r}^{0},\bar{x}^{j}_{s+h-1},\tilde{x}^{j}_{s+h-1},\bar{x}_{s}^{j+1},\tilde{x}_{s}^{j+1}$ (lines 4, 13, 19) queried for predictions stay within the feasible set $\mathcal{X}$.